# LLMBind: A Unified Modality-Task Integration Framework

### v3

@@
Abstract
While recent progress in multimodal large language models tackles various modality tasks, they posses limited integration capabilities for complex multi-modality tasks, consequently constraining the development of the field. In this work, we take the initiative to explore and propose the LLMBind, a unified framework for modality task integration, which binds Large Language Models and corresponding pre-trained task models with task-specific tokens. Consequently, LLMBind can interpret inputs and produce outputs in versatile combinations of image, text, video, and audio. Specifically, we introduce a Mixture-of-Experts technique to enable effective learning for different multimodal tasks through collaboration among diverse experts. Furthermore, we create a multi-task dataset comprising 400k instruction data, which unlocks the ability for interactive visual generation and editing tasks. Extensive experiments show the effectiveness of our framework across various tasks, including image, video, audio generation, image segmentation, and image editing. More encouragingly, our framework can be easily extended to other modality tasks, showcasing the promising potential of creating a unified AI agent for modeling universal modalities.
Machine Learning, ICML
1 Introduction
@@
Figure 1: Multimodal Large Language Models (MLLMs) are currently designed in three ways: (a) to obtain understanding ability by adding an encode-side adapter; (b) to achieve generation ability by decode-side adapter or (c) by text prompt. Differently, our method achieves modality-task integration by (d) task-specific text prompts and token embedding.
@@
Figure 2: Example Demonstrations. The LLMBind is designed to facilitate a user-friendly AI agent for achieving human-like conversation, interactive image, video, and audio generation, as well as interactive image editing and segmentation, among other tasks. Users can engage with LLMBind through multi-turn language dialogues, during which the model automatically identifies the relevant modality task and selects the appropriate model to accomplish the task.
@@
Recently, there has been significant progress in the field of AI, including the Large Language Models (LLMs) for natural language processing tasks and the diffusion models for high-resolution image generation. As shown in Figure 1, existing Multimodal Large Language Models (MLLMs) can be broadly categorized into three classes and we specifically use two metrics to evaluate the efficacy of a framework. The first indicator, called “updatability”, quantifies the ease of replacing the freezed task-specific model while the second, “versatility”, indicates the ability of the framework to extend to other tasks. The first class of methods, such as MiniGPT-4 (Zhu et al., 2023b) and LLaVA (Liu et al., 2023c), typically utilize encode-side adapters to handle inputs beyond language and can only perform the visual question-answering task. The second class of methods (Wu et al., 2023; Ghosal et al., 2023; Koh et al., 2023), as shown in Figure 1(b), incorporates decode-side adapters for various generation models, enhancing the adaptation of LLMs to complex multimodal tasks like generation. However, these methods lack updatability as they require training a new adapter when introducing a new generation model. For easy replacement of different generation models without additional training, the third class of methods (Xia et al., 2023; Brooks et al., 2023) directly generates desired text prompts for specific modality tasks, including image generation, audio generation, or image editing, as depicted in Figure 1(c). But these approaches can only perform tasks that only have text prompts as additional inputs, and cannot solve more complex visual tasks such as image segmentation. In brief, existing methods tend to solve limited tasks and do not explore the unification of more tasks within a shared LLM.
@@
In this work, we introduce LLMBind, a unified model integration framework designed to integrate models of various modality tasks, thereby unlocking the multitasking capabilities of the MLLM. As depicted in Figure 1(d), LLMBind is capable of processing two modalities: image and text. It processes these inputs through a visual encoder (Radford et al., 2021) and LLM (Chiang et al., 2023), ultimately producing task-specific text prompts or token embeddings. For tasks such as video generation (Ding et al., 2021b; Rombach et al., 2022a), image generation (Nichol et al., 2022; Tang et al., 2023), audio generation (Liu et al., 2023b; Yang et al., 2023a), and image editing (Brooks et al., 2023; Fu et al., 2023), we utilize text prompts as input to existing pre-trained generation models, significantly reducing the training complexity of our framework and enhancing its updateability. For modality tasks like image segmentation (Lai et al., 2023; Liu et al., 2023a), classification (Yang et al., 2023c; Naeem et al., 2023), and detection (Ge et al., 2021; Pi et al., 2023) that cannot be solely addressed with text prompts, we extract relevant semantic visual information from images using learnable task-specific token embeddings. Subsequently, these embeddings are utilized as an auxiliary input to the corresponding modality-task models.
@@
To guide the LLM to generate task-specific text prompts and token embeddings, we expand the vocabulary of LLM with learnable task-specific tokens, which can be divided into two categories: Task-Prompt Tokens and Semantic-Embedding Tokens. By training the LLM with specially designed instruction data pairs and task-specific tokens, we can instruct the model to discriminate the task type from user input and generate the corresponding text prompt or token embedding for a given multimodal task. Moreover, to better unify different modality tasks within the shared LLM, we introduce a highly effective technique called Mixture-of-Experts (MoE), which has been used successfully in recent studies (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022; Riquelme et al., 2021; Ahmed et al., 2016; Fedus et al., 2022). By incorporating this technique, our LLMBind can handle various modality tasks by utilizing different experts to collaboratively process multi-modal tasks, as shown in Figure 2.
Another fundamental contribution of this work is the creation of a dataset containing 400k instruction data for interactive generation and editing tasks. Each sample is generated by ChatGPT and formatted as a multi-round dialogue. This dataset addresses the scarcity of generation and editing instruction data and empowers LLMBind with robust interaction capabilities for generating and editing tasks.
Overall, LLMBind demonstrates promising results in advancing the development of human-like MLLM and AI agents. Contributions of this work are as follows:
  • We design a unified model integration framework that expands task-specific tokens for diverse modality tasks, thus easily integrating different tasks into a unified LLM, where we introduce the MoE technique in our framework to better handle diverse modality tasks.
  • We propose a unified MLLM that is compatible with various modality tasks, including image segmentation, image generation, image editing, video generation, and audio generation.
  • To facilitate the development of user-friendly interactive tasks, we construct a dataset of 400k interactive generation and editing multi-turn dialogues using ChatGPT. We plan to release this dataset as an open resource to foster collaborative advancements in this field.
Table 1: Comparison of functionalities among different MLLMs. indicates capability, indicates limitation, and “-” hints at the potential for easy extension. We cover a range of tasks including text conversation (Text-Con), image comprehension (Img-Com), image generation (Img-Gen), image editing (Img-Edit), video generation (Vid-Gen), audio generation (Aud-Gen), image segmentation (Img-Seg), and image detection (Img-Det).
MLLMs Text-Con Img-Com Img-Gen Img-Edit Vid-Gen Aud-Gen Img-Seg Img-Det LLaMA (Touvron et al., 2023) Vicuna (Chiang et al., 2023) Flaminggo (Alayrac et al., 2022) BLIP2 (Li et al., 2023b) mPLUG-Owl (Ye et al., 2023) DetGPT (Pi et al., 2023) LLaVA (Liu et al., 2023c) Mini-GPT4 (Zhu et al., 2023b) Audiogpt (Huang et al., 2023) Otter (Li et al., 2023a) LISA (Lai et al., 2023) GLaMM (Rasheed et al., 2023) NExTGPT (Wu et al., 2023) LLMBind (Ours) -
2 Related Work
2.1 Cross-Modal Understanding and Segmentation
With the introduction of frameworks such as LanguageBind (Zhu et al., 2023a) and ImageBind (Girdhar et al., 2023), which combine multi-modal data in a shared semantic space, several works focusing on cross-modal understanding (Liu et al., 2023c; Zhu et al., 2023b; Yin et al., 2023; Jin et al., 2023) have emerged. Along with these advancements, there has been a growing interest in text-guided image referring segmentation, a new approach that enables users to interact with images using user instruction. Unlike traditional semantic segmentation, which assigns class labels to pixels, referring segmentation aims to segment target objects based on complex textual instructions. Representative works in this area include LISA (Lai et al., 2023), GLaMM (Rasheed et al., 2023), GRES (Liu et al., 2023a), and LAVT (Yang et al., 2022).
2.2 Cross-Modal Generation and Editing
@@
Multi-modal generation has remained a highly active research area in recent years. These models (Ramesh et al., 2022; Rombach et al., 2022b; Hong et al., 2022b; Kong et al., 2020; Yang et al., 2023b) allow the generation of corresponding images, videos, and audio based on specified text prompts. However, this approach requires users to directly write text prompts for image generation, lacking interactive capabilities. Another approach to cross-modal generation (Wu et al., 2023; Zeqiang et al., 2023), combines dialogue with generation, making it more user-friendly and aligned with current user needs. Given that generated images may not always meet user requirements, image editing has emerged as another active research area within the field of artificial intelligence-generated content. Representative works in this area include Text2live (Bar-Tal et al., 2022), Imagic (Kawar et al., 2023), Sdedit (Meng et al., 2021).
2.3 Mixture of Experts
@@
The concept of Mixture of Experts (MoE) in artificial neural networks has evolved over time. Early methods activated all experts, leading to high computational requirements (Eigen et al., 2013). Subsequent research introduced sparse MoE models, where distinct experts are selectively activated using a learnable gating mechanism (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022). These approaches have shown success in computer vision (Eigen et al., 2013; Ahmed et al., 2016; Gross et al., 2017; Wang et al., 2020) and natural language processing (Shazeer et al., 2017; Lepikhin et al., 2020; Zhou et al., 2022) domains. Notably, the recent Mixtrial-8x7B model (mistralai, 2023) has garnered attention by showcasing the benefits of leveraging different experts for different tasks, resulting in improved performance while reducing computational costs.
2.4 Multimodal Large Language Models
@@
Recently, the emergence of Large Language Models (LLM) has ushered AI into a new era. With the open-sourcing of LLM models like LLaMA (Touvron et al., 2023) and Vicuna (Chiang et al., 2023), numerous research works have successfully explored methods to extend LLMs to other modalities, including video, audio, and images. This has led to the development of a variety of related MLLM models, as shown in Table 1. For instance, Flamingo (Alayrac et al., 2022) employs cross-attention structures for visual-text interaction, while DetGPT (Pi et al., 2023) connects multimodal LLMs with detectors, enabling object detection based on user instructions. BLIP-2 (Li et al., 2023b) encodes input images into queries using the Q-Former structure. mPLUG-Owl (Ye et al., 2023) and LLaVA (Liu et al., 2023c) encode images into image features using image encoders, which are then combined with text embeddings for LLM learning. LISA (Lai et al., 2023) achieves referring segmentation by training LLMs to output mask token embeddings in an end-to-end manner. NExTGPT (Wu et al., 2023) implements multimodal generation by aligning the output of the adapter, which is added after the LLM, with the output text features from the text encoder of various generation models (Liu et al., 2023b; Rombach et al., 2022a).
Figure 3: LLMBind architecture. Utilizing task-specific tokens, our framework implements universal multimodal understanding, segmentation, editing, and generation capabilities. During training, we employ LoRA-MoE techniques to facilitate model learning of different modality tasks and reduce computational complexity.
3 Method
@@
3.1 LLMBind
Overview. The framework of LLMBind is illustrated in Figure 3. It primarily consists of a vision transformer (Radford et al., 2021) as the image encoder, a large language model called Vicuna (Chiang et al., 2023), a SAM-like mask decoder as well as image, audio, and video diffusion models (Rombach et al., 2022a; Tang et al., 2023; Singer et al., 2022) for generation and editing tasks.
Visual Encoder and Tokenizer. We prioritize the processing of inputs from two primary modalities: images and text. For image encoding, we employ the ViT-L/14 (Radford et al., 2021) model to extract image features. We first reshape the image $x\in\mathbb{R}^{H\times W\times C}$ into a sequence of flattened 2D patches $X_{v}=[x_{v}^{1},x_{v}^{2},...,x_{v}^{N}]\in\mathbb{R}^{N\times(P^{2}\times C)}$, where $(H,W)$ is the shape of the original image, $C$ is the number of channels, $N=\frac{H\times W}{P^{2}}$ is the resulting number of image patches. The LLM uses constant latent vector size $D$ through all of its layers, so we use two MLP layers to map the image patches to $D$ dimensions. Similarly, text data is processed through a text embedding function $g$, which projects it into a sequence of text tokens $X_{t}=[x_{t}^{1},x_{t}^{2},\ldots,x_{t}^{N}]\in\mathbb{R}^{N\times D}$, where $N$ is the length of this text token sequence.
Task-Specific Tokens. Different multimodal tasks often require distinct network structures, making it difficult to consolidate them within a shared LLM. To overcome this challenge, we expand the vocabulary of the LLM with specially designed, learnable task-specific tokens. As shown in Figure 4 , these tokens can be classified into two categories: Task-Prompt Tokens and Semantic-Embedding Tokens. The first category includes tokens similar to standard text tokens, such as <gen> and </gen> for image generation, <vid_cap> and </vid_cap> for video generation, and <edit> and </edit> for image editing. These tokens appear in pairs, and the content between them represents the intermediate text prompt required for the corresponding modality-task. In contrast, the second category is imbued with dense semantic information, diverging from traditional text tokens. For example, <seg> contains rich visual information related to segmented objects in an image. After decoding by a mask decoder, this token can be converted back into masks. This category also encompasses tokens like <detect> and <cls>, designed for extracting spatial coordinates and classification data in image-based tasks. This dual-token framework not only enhances the adaptability of the LLM across various modalities but also ensures scalability and extensibility.
Figure 4: Task-Specific tokens include two categories: Task-Prompt Tokens and Semantic-Embedding Tokens.

### v4

@@
Abstract.
In the multi-modal domain, the dependence of various models on specific input formats leads to user confusion and hinders progress. To address this challenge, we introduce LLMBind, a novel framework designed to unify a diverse array of multi-modal tasks. By harnessing a Large Language Model (LLM) via Mixture-of-Experts (MoE), LLMBind processes multi-modal inputs and generates task-specific tokens, enabling the invocation of corresponding models to accomplish tasks. This unique approach empowers LLMBind to interpret inputs and generate outputs across various modalities, including image, text, video, and audio. Furthermore, we have constructed an interaction dataset comprising 400k instructions, which unlocks the ability of LLMBind for interactive visual generation and editing tasks. Extensive experimentation demonstrates that LLMBind achieves very superior performance across diverse tasks and outperforms existing models in user evaluations conducted in real-world scenarios. Moreover, the adaptability of LLMBind allows for seamless integration with the latest models and extension to new modality tasks, highlighting its potential to serve as a unified AI agent for modeling universal modalities. [^1] The source code will be made publicly available.
AI Agent, Multimodal Large Language Models, Mixture-of-Experts
ccs: Computing methodologies Multi-agent planningccs: Computing methodologies Computer vision tasksccs: Computing methodologies Robotic planning
1. Introduction
@@
Figure 1. Existing Multimodal Large Language Models are currently designed in three ways: (a) to obtain understanding ability by adding an encode-side adapter; (b) to achieve generation ability by decode-side adapter or (c) by text prompt. Differently, our method achieves modality-task integration by (d) task-specific text prompts and token embedding.
@@
Figure 2. Example Demonstrations. The LLMBind is designed to facilitate a user-friendly AI agent for achieving human-like conversation, interactive image, video, and audio generation, as well as interactive image editing and segmentation.
Recently, there has been significant progress in the field of Artificial Intelligence, including the Large Language Models (LLMs) (Chiang et al., 2023; Touvron et al., 2023) for natural language processing tasks and the diffusion models (Nichol et al., 2022; Tang et al., 2023; Liu et al., 2023a) for high-quality modality generation. Nevertheless, these models are typically limited to few modalities and tasks, highlighting the necessity to develop Multi-Modal Language Models (MLLMs) to integrate and address a broader range of tasks.
@@
Existing methods tend to solve limited tasks and do not explore the unification of more tasks within a shared LLM. As shown in Fig. 1, we first use two characteristics to evaluate the efficacy of a framework, the updateability which quantifies the ease of replacing the frozen task-specific model, and the expandability which indicates the ability of the framework to extend to other tasks. Then, we summarize the existing MLLMs into three categories: 1) Encode-side Adapter, such as MiniGPT-4 (Zhu et al., 2023a) and LLaVA (Liu et al., 2023c), typically utilize encode-side adapters to handle inputs beyond language and can only perform the visual question-answering task; 2) Decode-side Adapter (Wu et al., 2023; Ghosal et al., 2023; Koh et al., 2023), as shown in Fig. 1(b), incorporates decode-side adapters for various generation models, enhancing the adaptation of LLMs to complex multimodal tasks like generation. However, these methods lack updateability as they require training a new adapter when introducing a new generation model; 3) Text Prompt Agent (Xia et al., 2023; Brooks et al., 2023), as depicted in Fig. 1(c), can directly generate desired text prompts for specific modality tasks without additional training. However, these approaches can only perform tasks that only have text prompts as additional inputs, and cannot solve more complex visual tasks such as image segmentation.
@@
In this work, to achieve both updateability and expandability, we introduce LLMBind, a unified framework to integrate models for various modality tasks. As depicted in Fig. 1(d), LLMBind is capable of processing two modalities: image and text. It processes these inputs through a visual encoder (Radford et al., 2021) and LLM (Chiang et al., 2023), ultimately producing task-specific text prompts or token embeddings. For tasks such as video generation (Ding et al., 2021b; Rombach et al., 2022a), image generation (Nichol et al., 2022; Tang et al., 2023), audio generation (Liu et al., 2023a; Yang et al., 2023b), and image editing (Brooks et al., 2023; Fu et al., 2023), we utilize text prompts as input to existing pre-trained generation models, significantly reducing the training complexity of our framework and enhancing its updateability. For modality tasks like image segmentation (Lai et al., 2023; Liu et al., 2023b), classification (Yang et al., 2023a; Naeem et al., 2023), and detection (Ge et al., 2021) that cannot be solely addressed with text prompts, we extract relevant semantic visual information from images using learnable task-specific token embeddings. Subsequently, these embeddings are utilized as an auxiliary input to the corresponding modality-task models.
@@
To guide the LLM to generate task-specific text prompts and token embeddings, we expand the vocabulary of LLM with learnable task-specific tokens, which can be divided into two categories: Task-Prompt Tokens and Semantic-Embedding Tokens. By training the LLM with specially designed instruction data pairs and task-specific tokens, we can instruct the model to discriminate the task type from user input and generate the corresponding text prompt or token embedding for a given multimodal task. Moreover, to better unify different modality tasks within the shared LLM, we introduce a highly effective technique called Mixture-of-Experts (MoE), which has been used successfully in recent studies (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022; Riquelme et al., 2021; Ahmed et al., 2016; Fedus et al., 2022). By incorporating this technique, our LLMBind can handle various modality tasks by utilizing different experts to collaboratively process multi-modal tasks, as shown in Fig. 2.
In practical applications, users often require multi-turn interactive generation and editing to achieve precise and satisfactory results. For example, in image editing, users need to apply various effects repeatedly to achieve their desired visual outcome. To this end, we develop a dataset comprising 400k instructional samples for interactive generation and editing tasks. Each sample, generated by ChatGPT, is formatted as a multi-round dialogue. This dataset alleviates the scarcity of instruction data for generation and editing tasks, equipping LLMBind with robust interaction capabilities.
Extensive experiments have demonstrated LLMBind’s superior performance in tasks such as referring segmentation, video generation, image generation, and audio generation, surpassing numerous specialized and multitask models. Additionally, LLMBind has exhibited remarkable performance in human evaluations of interactive generation tasks, exceeding traditional generation methods in terms of both the quality of generated data and its alignment with text. Overall, LLMBind demonstrates promising results in advancing the development of human-like MLLM and AI agents. Our Contributions to this work are as follows:
  • We introduce LLMBind, a unified framework for various modality tasks, encompassing image segmentation, generation, editing, as well as video and audio generation. We employ the MoE technique to effectively handle diverse tasks.
  • We devise a task-specific token strategy to seamlessly integrate different tasks into a unified MLLM, ensuring high model’s updateability and expandability for emerging tasks.
  • We curate an interactive dataset comprising 400k instructional data, enabling LLMBind to tackle multi-turn interactive visual generation and editing tasks.
2. Related Work
2.1. Cross-Modal Generation and Editing
@@
Multi-modal generation has remained a highly active research area in recent years. These models (Ramesh et al., 2022; Rombach et al., 2022b; Hong et al., 2022a; Kong et al., 2020; Yang et al., 2023c) allow the generation of corresponding images, videos, and audio based on specified text prompts. However, this approach requires users to directly write text prompts for image generation, lacking interactive capabilities. Another approach to cross-modal generation (Wu et al., 2023; Zeqiang et al., 2023), combines dialogue with generation, making it more user-friendly and aligned with current user needs. Given that generated images may not always meet user requirements, image editing has emerged as another active research area within the field of artificial intelligence-generated content. Representative works in this area include Text2live (Bar-Tal et al., 2022), Imagic (Kawar et al., 2023), Sdedit (Meng et al., 2021).
2.2. Cross-Modal Segmentation
With the introduction of frameworks such as LanguageBind (Zhu et al., 2023b) and ImageBind (Girdhar et al., 2023), which combine multi-modal data in a shared semantic space, several works focusing on cross-modal understanding (Liu et al., 2023c; Zhu et al., 2023a; Yin et al., 2023; Jin et al., 2023) have emerged. Along with these advancements, there has been a growing interest in text-guided image referring segmentation, a new approach that enables users to interact with images using user instruction. Unlike traditional semantic segmentation, which assigns class labels to pixels, referring segmentation (Lai et al., 2023; Liu et al., 2023b) aims to segment target objects based on complex textual instructions.
Table 1. Comparison of functionalities among different MLLMs. indicates capability, indicates limitation, and “-” hints at the potential for easy extension. We cover a range of tasks including text conversation (Text-Con), image comprehension (Img-Com), image generation (Img-Gen), image editing (Img-Edit), video generation (Vid-Gen), audio generation (Aud-Gen), image segmentation (Img-Seg), and image detection (Img-Det).
MLLMs Text-Con Img-Com Img-Gen Img-Edit Vid-Gen Aud-Gen Img-Seg Img-Det LLaMA (Touvron et al., 2023) Vicuna (Chiang et al., 2023) Flaminggo (Alayrac et al., 2022) BLIP2 (Li et al., 2023a) mPLUG-Owl (Ye et al., 2023) DetGPT (Pi et al., 2023) LLaVA (Liu et al., 2023c) Mini-GPT4 (Zhu et al., 2023a) Audiogpt (Huang et al., 2023) Otter (Li et al., 2023b) LISA (Lai et al., 2023) GLaMM (Rasheed et al., 2023) NExTGPT (Wu et al., 2023) LLMBind (Ours) -
2.3. Mixture-of-Experts
@@
The concept of Mixture-of-Experts (MoE) in artificial neural networks has evolved. Early methods activated all experts, leading to high computational requirements (Eigen et al., 2013). Subsequent research introduced sparse MoE models, where distinct experts are selectively activated using a learnable gating mechanism (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022). These approaches have shown success in computer vision (Eigen et al., 2013; Ahmed et al., 2016; Gross et al., 2017; Wang et al., 2020) and natural language processing (Shazeer et al., 2017; Lepikhin et al., 2020; Zhou et al., 2022) domains.
2.4. Multimodal Large Language Models
@@
Recently, the emergence of Large Language Models (LLM) has ushered AI into a new era. With the open-sourcing of LLM models like LLaMA (Touvron et al., 2023) and Vicuna (Chiang et al., 2023), numerous research works have successfully explored methods to extend LLMs to other modalities, including video, audio, and images. This has led to the development of a variety of related MLLM models, as shown in Table 1. For instance, Flamingo (Alayrac et al., 2022) employs cross-attention structures for visual-text interaction, while DetGPT (Pi et al., 2023) connects multimodal LLMs with detectors, enabling object detection based on user instructions. BLIP-2 (Li et al., 2023a) encodes input images into queries using the Q-Former structure. mPLUG-Owl (Ye et al., 2023) and LLaVA (Liu et al., 2023c) encode images into image features using image encoders, which are then combined with text embeddings for LLM learning. LISA (Lai et al., 2023) achieves referring segmentation by training LLMs to output mask token embeddings in an end-to-end manner. NExTGPT (Wu et al., 2023) implements multimodal generation by aligning the output of the adapter, which is added after the LLM, with the output text features from the text encoder of various generation models (Liu et al., 2023a; Rombach et al., 2022a).
3. Methodology
@@
3.1. LLMBind
Overview. The framework of LLMBind is illustrated in Fig. 3. It primarily consists of a vision transformer (Radford et al., 2021) as the image encoder, a large language model called Vicuna (Chiang et al., 2023), a SAM-like mask decoder as well as image, audio, and video diffusion models (Rombach et al., 2022a; Tang et al., 2023; Singer et al., 2022b) for generation and editing tasks.
Figure 3. LLMBind architecture. Utilizing task-specific tokens, our framework implements universal multimodal understanding, segmentation, editing, and generation capabilities. During training, we employ LoRA-MoE techniques to facilitate model learning of different modality tasks and reduce computational complexity.
Visual Encoder and Tokenizer. We prioritize the processing of inputs from two primary modalities: images and text. For image encoding, we employ the ViT-L/14 (Radford et al., 2021) to extract image features. We first reshape the image $x\in\mathbb{R}^{H\times W\times C}$ into a sequence of flattened 2D patches $X_{v}=[x_{v}^{1},x_{v}^{2},...,x_{v}^{N}]\in\mathbb{R}^{N\times(P^{2}\times C)}$, where $(H,W)$ is the shape of the original image, $C$ is the number of channels, $N=\frac{H\times W}{P^{2}}$ is the resulting number of image patches. The LLM uses constant latent vector size $D$ through all of its layers, so we use two MLP layers to map the image patches to $D$ dimensions. Similarly, text is processed through a text embedding function $g$, which projects it into a sequence of text tokens $X_{t}=[x_{t}^{1},x_{t}^{2},\ldots,x_{t}^{N}]\in\mathbb{R}^{N\times D}$, where $N$ is the length of this text token sequence.
Task-Specific Tokens. Different multimodal tasks often require distinct network structures, making it difficult to consolidate them within a shared LLM. To overcome this challenge, we expand the vocabulary of the LLM with specially designed, learnable task-specific tokens. As shown in Fig. 4 , these tokens can be classified into two categories: Task-Prompt Tokens and Semantic-Embedding Tokens. The first category includes tokens similar to standard text tokens, such as <gen> and </gen> for image generation, <vid_cap> and </vid_cap> for video generation, and <edit> and </edit> for image editing. These tokens appear in pairs, and the content between them represents the intermediate text prompt required for the corresponding modality task. In contrast, the second category is imbued with dense semantic information, diverging from traditional text tokens. For example, <seg> contains rich visual information related to segmented objects in an image. After decoding by a mask decoder, this token can be converted back into masks. This category also encompasses tokens like <detect> and <cls>, designed for extracting spatial coordinates and classification data in image-based tasks. This dual-token framework not only enhances the adaptability of the LLM across various modalities but also ensures scalability and extensibility.
Figure 4. Task-Specific tokens include two categories: Task-Prompt Tokens and Semantic-Embedding Tokens.